OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 50 km/h. After six hours, the velocity of the car is 54 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. Part B: How can you graph the equation obtained in Part A for the first six hours?

OpenStudy (ageta):

that means (2, 54) and (4, 58) are two points on the required graph and u can use these to write the equation for velocity in slope-intercept form

OpenStudy (ageta):

start by finding the slope

OpenStudy (ageta):

(2,54) (4,58) 58-54=4/2= 2

OpenStudy (ageta):

Is it v=2t+50?

OpenStudy (anonymous):

Wouldn't the point be (6,54) because at six hours it's 54 km/h ?

OpenStudy (ageta):

yes, plugging in t = 6 gives u the endpoint ^

OpenStudy (ageta):

Your equation for velocity from part A : v=2t+50 when t=0, v=? when t=6, v=?

OpenStudy (anonymous):

When t=0 then v=50 and when t=6 then v=62

OpenStudy (ageta):

yes

OpenStudy (anonymous):

So my equation for part a would be v=2t+50 ? But now how do we graph it

OpenStudy (ageta):

yes

OpenStudy (anonymous):

Oh and I made a mistake for part b it's suppose to say at the end for the first seven hours not six

OpenStudy (ageta):



OpenStudy (ageta):

if u draw a vertical on any point along the x axis... it will never cut twice at the same point

OpenStudy (ageta):